OpenStudy (anonymous):

With a hypotenuse of 6, what is the length of one leg?

OpenStudy (anonymous):

in a 45-45-90 triangle. The legs are equal

OpenStudy (anonymous):

if it is iscoceles then 36=2x^2 x^2=18 x= sqrt18

Parth (parthkohli):

Thus,\[a^2 + b^2 = 6^2 \\ a^2 + a^2 = 6^2\]

OpenStudy (anonymous):

\[(6)^2=(x)^2+(x)^2=2x^2=36\]

OpenStudy (anonymous):

i got 18 but here is the multiple choice that is confusing me. A 11sqrt3 B4sqrt7 C7sqrt8 D3sqrt2 E 8sqrt9 F 2sqrt5

OpenStudy (anonymous):

\[\sqrt{18}=\sqrt{3*3*2}=3\sqrt2\]

OpenStudy (anonymous):

@PetraRenee1212 does that help?

OpenStudy (anonymous):

oh wow your good thanks alot. :D

OpenStudy (anonymous):

:)YW